# Adapting Noise to Data: Generative Flows from 1D Processes

### v2

@@ 4.2 Learning Quantile Processes
where $\pi_{\phi}\in\Pi_{o}(\mu_{0},\nu_{\phi})$, see Remark 2. In practice we optimize empirical expectations over minibatches; we compute a mini-batch OT coupling once per batch and use it within both loss terms, see Appendix D.4.
@@ 5 Experiments
To provide intuition and validate our proposed method, we conduct experiments on both synthetic and image datasets. For each component, we model the quantile with a Rational Quadratic Spline (RQS) [14, 12] and add a learnable scale and bias. This keeps monotonicity, is parameter-efficient, and gives analytic derivatives. See Appendix D.2 for details. The code is available on GitHub at https://github.com/TUB-Angewandte-Mathematik/Adapting-Noise.
@@ 5.1 Synthetic Datasets
We begin by qualitatively analyzing our algorithm on several synthetic 2D distributions (see Appendix D.1), each designed to highlight a specific aspect of our approach. We provide intuition about the learned latent distribution and demonstrate that it is closer to the data in the Wasserstein sense, yields shorter transport paths, and successfully captures the tail behavior.
@@
The funnel distribution, shown in Figure 5, presents a challenge due to its heavy-tailed, conditional structure. Several methods for handling heavy-tailed datasets have already been introduced in the context of diffusion models [25, 30, 31]. For example, we compare our funnel example to [25], where the authors hand-select the parameters of a Student-$t$ distribution in each dimension.
@@ 5.2 Image Datasets
In high-dimensional settings and given fixed batch sizes, the signal for the quantile function can be noisy, potentially leading to degenerate solutions. To mitigate this, we add a regularization term to the loss that penalizes the expected negative log-determinant of the Jacobian of the quantile. Access to analytic derivatives makes this computation efficient, for more details see D.3.
@@
In Figure  11, we compare the learned and empirical quantiles on the MNIST dataset at different pixel locations $(x,y)$. Where the pixel is essentially black, the learned quantile concentrates around that value, whereas in the center regions, where uncertainty is higher, the quantiles remain spread around zero (gray), accurately reflecting the data variability.
@@
In Figure  12, we compare the performance under different network capacity constraints by evaluating our learned latent against a Gaussian latent. Both latents are trained using mini-batch optimal transport. As already observed in Fig. 10, the transport paths are significantly shorter since the learned latent successfully minimizes the Wasserstein distance and removes redundant information. This enables the network to use the available parameters more efficiently and achieve better results with the same parameter count.
@@ Appendix A Uniform Process and MMD Gradient Flow
and $Q_{\nu}(s)=a(1-s)+bs$. it was shown in [17] that the functional $F_{\nu}\colon L_{2}(0,1)\to\mathbb{R}$ defined by
@@
fulfills $\mathcal{F}_{\nu}(\mu)=F_{\nu}(Q_{\mu})$ for all $\mu\in\mathcal{P}_{2}(\mathbb{R})$. Moreover, we have the following equivalent characterization of Wasserstein gradient flows of $\mathcal{F}_{\nu}$, which can be found in [11, Theorem 4.5].
@@ B.2 Uniform Case
Finally we want to note that the MMD functional (27) loses its convexity (along generalized geodesics) in multiple dimensions [17], and the general existence of its Wasserstein gradient flows is unclear in the multivariate case. This yields another reason to work in 1D.
@@ Appendix C IMM with Quantile Interpolants
In this section, we want to demonstrate how the IMM framework proposed in [41] can be realized by our quantile approach. Note that in the following – for notational simplicity – we consider the one-dimensional case $X_{0},Z_{t}\in\mathbb{R}$ where we can employ quantile functions. By combining the 1D components into a multivariate model $\mathbf{X}_{0}=(X_{0}^{1},...,X_{0}^{d}),\,\mathbf{Z}_{t}=(Z_{t}^{1},...,Z_{t}^{d})$, the results of this chapter trivially extend to $\mathbb{R}^{d}$.
@@
The interpolants used in Denoising Diffusion Implicit Models (DDIMs) [33] are given by
@@
Proposition 12 represents the key observation which allows us to utilize our quantile process (63) in the IMM framework the same way as [41] employ the DDIM interpolants (65):
@@
The following proposition follows directly from Proposition 12 as in [41]. It is essential for deriving the appropriate loss functions.
@@
Then it is proposed in [41, Eq. (7)] to minimize the so-called naïve objective
@@
Although for marginal-preserving interpolants, a minimizer of $\mathcal{L}_{\rm naive}$ exists with minimum 0, the authors of [41] object that directly optimizing (78) faces practical difficulties when $t$ is far away from $s$. Instead, they propose to apply the following “inductive bootstrapping” technique:
@@
Thus, for the number of steps $n\to\infty$, it holds $0={\rm MMD}^{2}_{K}(\rho_{s},p^{\theta_{n}}(s,t_{n}))$ even for the entire interval $t_{n}\in[s,1]$. Hence, minimizing the general objective (79) with a large number of steps eventually minimizes the naïve objective (78), see [41, Theorem 1] for more details.
Appendix D Adapting Noise to Data
D.1 Toy Target Distributions
@@
For the toy illustration in Figure 5, we work with the dataset known as Neals Funnel [24]. The distribution of Neal’s funnel is defined as follows:
@@
D.2 Details on the Architecture of the Learned Quantiles
We implement each one–dimensional quantile function with rational–quadratic splines (RQS) [14, 12]. We explored several ways to map $u\in(0,1)$ into the spline input; the two variants below consistently performed well and are used in our experiments. For every coordinate $i$, we write
@@ D.2 Details on the Architecture of the Learned Quantiles
D.3 Regularization via Expected Negative Log–Jacobian
@@
D.4 Minibatch Optimal Transport
@@
Appendix E Implementation Details
@@
The coupling plans are calculated using the Python Optimal Transport package [13]. For inference simulate the corresponding ODEs using the torchdiffeq [7] package. For all models we only used the batch size $128$ and learning rate 2e42e-4 for the velocities. Quantile transports are parameterised by stacked rational-quadratic splines as described in D.2, we set the minimum bin width and height to 1e31e-3 and the minimum slope to 1e51e-5.
E.1 Synthetic Examples
@@
The quantiles were trained for the first 20,000 steps, after which the learning rate was linearly decayed to 0 by step 25,000. For both datasets, we trained the velocity model with 4 layers and a hidden width of 256 for 100,000 steps. For the RQS we choose the parameters number of bins 32, bound 5, layers 3.
@@
E.2 Image Experiments
For both image datasets, we adapt the U-Net from [9] to parametrize our velocity field.
@@ E.2 Image Experiments
Here we use the U-Net with base width 128, channel multipliers $(1,2,2,2)$, two residual blocks per resolution, attention at $16\times 16$, four attention heads , and dropout 0.1. We clip the gradient norm to 1 and use exponential moving averaging with a decay of $0.9999$. To evaluate our results, we use the Fréchet inception distance (FID) [18]. The quantiles were trained for the first 20,000 steps, after which the learning rate was linearly decayed to 0 by step 25,000.

### v3

@@ 4.2 Learning Quantile Processes
where $\pi_{\phi}\in\Pi_{o}(\mu_{0},\nu_{\phi})$, see Remark 2. In practice we optimize empirical expectations over minibatches; we compute a mini-batch OT coupling once per batch and use it within both loss terms, see Appendix E.4.
@@ 5 Experiments
To provide intuition and validate our proposed method, we conduct experiments on both synthetic and image datasets. For each component, we model the quantile with a Rational Quadratic Spline (RQS) [GD1982, durkan2019neural] and add a learnable scale and bias. This keeps monotonicity, is parameter-efficient, and gives analytic derivatives. See Appendix E.2 for details. The code is available on GitHub at https://github.com/TUB-Angewandte-Mathematik/Adapting-Noise.
@@ 5.1 Synthetic Datasets
We begin by qualitatively analyzing our algorithm on several synthetic 2D distributions (see Appendix E.1), each designed to highlight a specific aspect of our approach. We provide intuition about the learned latent distribution and demonstrate that it is closer to the data in the Wasserstein sense, yields shorter transport paths, and successfully captures the tail behavior.
@@
The funnel distribution, shown in Figure 5, presents a challenge due to its heavy-tailed, conditional structure. Several methods for handling heavy-tailed datasets have already been introduced in the context of diffusion models [pandey2024heavytaileddiffusionmodels, shariatian2025heavytaileddiffusiondenoisinglevy, HTDL2025]. For example, we compare our funnel example to [pandey2024heavytaileddiffusionmodels], where the authors hand-select the parameters of a Student-$t$ distribution in each dimension.
@@ 5.2 Image Datasets
In high-dimensional settings and given fixed batch sizes, the signal for the quantile function can be noisy, potentially leading to degenerate solutions. To mitigate this, we add a regularization term to the loss that penalizes the expected negative log-determinant of the Jacobian of the quantile. Access to analytic derivatives makes this computation efficient, for more details see E.3.
@@
In Figure  12, we compare the learned and empirical quantiles on the MNIST dataset at different pixel locations $(x,y)$. Where the pixel is essentially black, the learned quantile concentrates around that value, whereas in the center regions, where uncertainty is higher, the quantiles remain spread around zero (gray), accurately reflecting the data variability.
@@
In Figure  13, we compare the performance under different network capacity constraints by evaluating our learned latent against a Gaussian latent. Both latents are trained using mini-batch optimal transport. As already observed in Fig. 11, the transport paths are significantly shorter since the learned latent successfully minimizes the Wasserstein distance and removes redundant information. This enables the network to use the available parameters more efficiently and achieve better results with the same parameter count.
@@ Appendix A Uniform Process and MMD Gradient Flow
and $Q_{\nu}(s)=a(1-s)+bs$. it was shown in [HGBS2022] that the functional $F_{\nu}\colon L_{2}(0,1)\to\mathbb{R}$ defined by
@@
fulfills $\mathcal{F}_{\nu}(\mu)=F_{\nu}(Q_{\mu})$ for all $\mu\in\mathcal{P}_{2}(\mathbb{R})$. Moreover, we have the following equivalent characterization of Wasserstein gradient flows of $\mathcal{F}_{\nu}$, which can be found in [DSBHS2024, Theorem 4.5].
@@ B.2 Uniform Case
Finally we want to note that the MMD functional (27) loses its convexity (along generalized geodesics) in multiple dimensions [HGBS2022], and the general existence of its Wasserstein gradient flows is unclear in the multivariate case. This yields another reason to work in 1D.
@@ Appendix C IMM with Quantile Interpolants
In this section, we want to demonstrate how the IMM framework proposed in [IMM2025] can be realized by our quantile approach. Note that in the following – for notational simplicity – we consider the one-dimensional case $X_{0},Z_{t}\in\mathbb{R}$ where we can employ quantile functions. By combining the 1D components into a multivariate model $\mathbf{X}_{0}=(X_{0}^{1},...,X_{0}^{d}),\,\mathbf{Z}_{t}=(Z_{t}^{1},...,Z_{t}^{d})$, the results of this chapter trivially extend to $\mathbb{R}^{d}$.
@@
The interpolants used in Denoising Diffusion Implicit Models (DDIMs) [DDIM2020] are given by
@@
Proposition 12 represents the key observation which allows us to utilize our quantile process (63) in the IMM framework the same way as [IMM2025] employ the DDIM interpolants (65):
@@
The following proposition follows directly from Proposition 12 as in [IMM2025]. It is essential for deriving the appropriate loss functions.
@@
Then it is proposed in [IMM2025, Eq. (7)] to minimize the so-called naïve objective
@@
Although for marginal-preserving interpolants, a minimizer of $\mathcal{L}_{\rm naive}$ exists with minimum 0, the authors of [IMM2025] object that directly optimizing (78) faces practical difficulties when $t$ is far away from $s$. Instead, they propose to apply the following “inductive bootstrapping” technique:
@@
Thus, for the number of steps $n\to\infty$, it holds $0={\rm MMD}^{2}_{K}(\rho_{s},p^{\theta_{n}}(s,t_{n}))$ even for the entire interval $t_{n}\in[s,1]$. Hence, minimizing the general objective (79) with a large number of steps eventually minimizes the naïve objective (78), see [IMM2025, Theorem 1] for more details.
Appendix D Radially Symmetric Processes
While our main construction relies on independent one-dimensional processes across coordinates (Section 3), we can alternatively build multi-dimensional processes using radial symmetry. This approach, sometimes called polar factorization, decomposes a random vector as the product of its radius and direction.
Let $x\in\mathbb{R}^{d}$, write $r=\|x\|$ and $\omega=x/\|x\|\in\mathbb{S}^{d-1}$. A vector field is called radial if
$v_{t}(x)=v_{t}^{R}(r)\,\omega,$ (82)
where $v_{t}^{R}:[0,\infty)\to\mathbb{R}$ is its radial component. Likewise, a time-dependent probability density $\rho_{t}\colon\mathbb{R}^{d}\to\mathbb{R}$ is radially symmetric if there exists $\bar{\rho}_{t}\colon[0,\infty)\to\mathbb{R}$ such that
$\rho_{t}(x)=\bar{\rho}_{t}\bigl(\|x\|\bigr)\quad\text{for all }x\in\mathbb{R}^{d}.$ (83)
Define the radial mass density
$f_{R}(t,r):=\int_{\mathbb{S}^{d-1}}\rho_{t}(r\omega)\,r^{d-1}\,\mathrm{d}\sigma(\omega)=|\mathbb{S}^{d-1}|\,r^{d-1}\,\bar{\rho}_{t}(r),$ (84)
where $\sigma$ denotes the surface measure on $\mathbb{S}^{d-1}$. If $\mathbf{X}\in\mathbb{R}^{d}$ has radially symmetric density $\rho$, then $f_{R}$ is precisely the probability density of the radius $\|\mathbf{X}\|$.
Lemma 14.
Let $\rho_{t}$ be a smooth, radially symmetric density on $\mathbb{R}^{d}$, and let
$v_{t}(x)=v_{t}^{R}(r)\,\frac{x}{r}$ (85)
be a smooth radial velocity field. Then, the pair $(\rho_{t},v_{t})$ satisfies the $d$-dimensional continuity equation (1) if and only if $(f_{R}(t,\cdot),v_{t}^{R})$ fulfills the one-dimensional flux equation for $r>0$:
$\partial_{t}f_{R}(t,r)+\partial_{r}\bigl[f_{R}(t,r)\,v_{t}^{R}(r)\bigr]=0,$ (86)
where $f_{R}$ is defined in (84).
Proof.
We compute the divergence in Cartesian coordinates:
$\nabla_{x}\!\cdot\bigl[\bar{\rho}_{t}(r)\,v_{t}^{R}(r)\,\tfrac{x}{r}\bigr]=\sum_{i=1}^{d}\partial_{x_{i}}\Bigl(\bar{\rho}_{t}(r)\,v_{t}^{R}(r)\,\tfrac{x_{i}}{r}\Bigr).$ (87)
Since $\bar{\rho}_{t}$ and $v_{t}^{R}$ depend only on $r=\|x\|$, and $\partial_{x_{i}}r=x_{i}/r$, the product rule gives
$\partial_{x_{i}}\Bigl(\bar{\rho}_{t}\,v_{t}^{R}\,\tfrac{x_{i}}{r}\Bigr)=\partial_{r}(\bar{\rho}_{t}\,v_{t}^{R})\,\frac{x_{i}^{2}}{r^{2}}+\bar{\rho}_{t}\,v_{t}^{R}\,\frac{1}{r}-\bar{\rho}_{t}\,v_{t}^{R}\,\frac{x_{i}^{2}}{r^{3}}.$ (88)
Summing over $i$ and using $\sum_{i=1}^{d}x_{i}^{2}=r^{2}$ yields
$\nabla_{x}\!\cdot(\bar{\rho}_{t}\,v_{t})=\partial_{r}(\bar{\rho}_{t}\,v_{t}^{R})+\frac{d-1}{r}\,\bar{\rho}_{t}\,v_{t}^{R}=r^{-(d-1)}\,\partial_{r}\bigl(r^{d-1}\,\bar{\rho}_{t}\,v_{t}^{R}\bigr).$ (89)
Hence the continuity equation (1) is equivalent to
$\partial_{t}\bar{\rho}_{t}+\frac{1}{r^{d-1}}\,\partial_{r}\bigl(r^{d-1}\bar{\rho}_{t}\,v_{t}^{R}\bigr)=0.$ (90)
Multiplying by $|\mathbb{S}^{d-1}|\,r^{d-1}$ shows the equivalence to (86). ∎
Definition 15 (Radial Processes).
Let $(R_{t})_{t}$ be a non-negative one-dimensional stochastic process with $R_{0}=0$, probability density $p_{t}^{R}$, and velocity field $v_{t}^{R}$ satisfying the one-dimensional continuity equation (13). Let $\mathbf{U}\sim\mathrm{Unif}(\mathbb{S}^{d-1})$ be uniformly distributed on the unit sphere, independent of $R_{t}$. The radial process w.r.t. $(R_{t})_{t}$ in $\mathbb{R}^{d}$ is defined by the polar factorization
$\mathbf{Y}_{t}=R_{t}\,\mathbf{U}.$ (91)
Since $\|\mathbf{Y}_{t}\|=|R_{t}|$, we have $f_{R}(t,r)=p_{t}^{R}(r)$ for all $r>0$.
The denotation of the velocity field of the process $(R_{t})_{t}$ as $v_{t}^{R}$ is justified by the following proposition.
Proposition 16 (Radial Continuity).
Let $\mu_{t}$ denote the law of $\mathbf{Y}_{t}$ in Definition 15, with radially symmetric density $\rho_{t}(x)=\bar{\rho}_{t}(\|x\|)$. Let $p_{t}^{R}$ and $v_{t}^{R}$ be given as in Definition 15.
Then, the velocity field
$v_{t}(x)\coloneqq v_{t}^{R}(\|x\|)\,\frac{x}{\|x\|}$ (92)
satisfies the $d$-dimensional continuity equation (1) together with $\mu_{t}$. In other words, the radial process $\mathbf{Y}_{t}$ admits a radial velocity field $v_{t}$ with radial component $v_{t}^{R}$ given by $R_{t}$.
Proof.
By Lemma 14, it suffices to verify (86). But we have
$\partial_{t}f_{R}(t,r)+\partial_{r}[f_{R}(t,r)\,v_{t}^{R}(r)]=\bigl[\partial_{t}p_{t}^{R}(r)+\partial_{r}(p_{t}^{R}\,v_{t}^{R})(r)\bigr]=0,$ (93)
since $(p_{t}^{R},v_{t}^{R})$ satisfies the one-dimensional continuity equation (13) by definition. ∎
Remark 17 (Connection to Mean-Reverting Processes and Learning).
The radial construction integrates with the mean-reverting framework from Section 2.3. Given $\mathbf{X}_{0}\sim\mu_{0}$ and scheduling functions $f,g$ satisfying (6), we define
$\mathbf{X}_{t}=f(t)\,\mathbf{X}_{0}+R_{g(t)}\,\mathbf{U},$ (94)
where $R_{t}$ is a radial process and $\mathbf{U}\sim\mathrm{Unif}(\mathbb{S}^{d-1})$ is independent of $\mathbf{X}_{0}$. For instance, using the Kac process as the underlying one-dimensional radial process yields a radially symmetric alternative to the componentwise construction.
For the scaled latent case, choosing f(t)=1tf(t)=1-t, $g(t)=t$, and $R_{t}=t\,R$ where $R\sim\chi_{d}$ follows the chi distribution with $d$ degrees of freedom (i.e., $R=\|\mathbf{Z}\|$ for $\mathbf{Z}\sim\mathcal{N}(0,I_{d})$), we recover the standard Gaussian latent $\mathbf{Y}_{1}\sim\mathcal{N}(0,I_{d})$.
The radial construction naturally extends to the quantile learning framework of Section 4. We can parametrize the radial distribution via its quantile function $Q_{R}:(0,1)\to[0,\infty)$, where $R_{t}=t\,Q_{R}(U)$ with $U\sim\mathcal{U}(0,1)$. Following Section 4.2, we learn $Q_{R}$, which can also be combined with velocity field training analogously to (41). This provides a data-adapted isotropic latent, in contrast to the fully independent coordinate-wise adaptation of our main method.
Appendix E Adapting Noise to Data
E.1 Toy Target Distributions
@@
For the toy illustration in Figure 5, we work with the dataset known as Neals Funnel [neal2003slice]. The distribution of Neal’s funnel is defined as follows:
@@
E.2 Details on the Architecture of the Learned Quantiles
We implement each one–dimensional quantile function with rational–quadratic splines (RQS) [GD1982, durkan2019neural]. We explored several ways to map $u\in(0,1)$ into the spline input; the two variants below consistently performed well and are used in our experiments. For every coordinate $i$, we write
@@ E.2 Details on the Architecture of the Learned Quantiles
E.3 Regularization via Expected Negative Log–Jacobian
@@
E.4 Minibatch Optimal Transport
@@
Appendix F Implementation Details
@@
The coupling plans are calculated using the Python Optimal Transport package [flamary2021pot]. For inference simulate the corresponding ODEs using the torchdiffeq [torchdiffeq] package. For all models we only used the batch size $128$ and learning rate 2e42e-4 for the velocities. Quantile transports are parameterised by stacked rational-quadratic splines as described in E.2, we set the minimum bin width and height to 1e31e-3 and the minimum slope to 1e51e-5.
F.1 Synthetic Examples
@@
The quantiles were trained for the first 20,000 steps, after which the learning rate was linearly decayed to 0 by step 25,000. For both datasets, we trained the velocity model with 3 layers and a hidden width of 64 for 100,000 steps. For the RQS we choose the parameters number of bins 32, bound 50, layers 3.
@@
F.2 Image Experiments
For both image datasets, we adapt the U-Net from [dhariwal2021diffusion] to parametrize our velocity field.
@@ F.2 Image Experiments
Here we use the U-Net with base width 128, channel multipliers $(1,2,2,2)$, two residual blocks per resolution, attention at $16\times 16$, four attention heads , and dropout 0.1. We clip the gradient norm to 1 and use exponential moving averaging with a decay of $0.9999$. To evaluate our results, we use the Fréchet inception distance (FID) [HRUNH2017]. The quantiles were trained for the first 20,000 steps, after which the learning rate was linearly decayed to 0 by step 25,000.